OpenStudy (anonymous):

A nonconducting spherical shell, with an inner radius of 4 cm and an outer radius of 6 cm, has charge spread non uniformly through its volume between its inner and outer surfaces. The volume charge density ρ is the charge per unit volume, with the unit coulomb per cubic meter. For this shell ρ = b/r where r is the distance in meters from the center of the shell and b = 3 μ C/m2. What is the net charge in the shell?

OpenStudy (anonymous):

dQ/dV = ρ In other words a infinitesimal segment dV caries the charge dQ = ρ dV Let dV be a spherical shell between between r and (r + dr): dV = (4π/3)·( (r + dr)² - r³ ) = (4π/3)·( r³ + 3·r²·dr + 3·r·(dr)² + /dr)³ - r³ ) = (4π/3)·( 3·r²·dr + 3·r·(dr)² + /dr)³ ) drop higher order terms = 4·π·r²·dr To get total charge integrate over the whole volume of your object, i.e. from ri to ra: Q = ∫ dQ = ∫ ρ dV = ∫ri→ra { (b/r)·4·π·r² } dr = ∫ri→ra { 4·π·b·r } dr = 2·π·b·( ra² - ri² ) With given parameters: Q = 2·π · 3µC/m²·( (6cm)² - (4cm)² ) = 2·π · 3×10⁻⁶C/m²·( (6×10⁻²m)² - (4×10⁻²m)² ) = 3.77×10⁻⁸C = 37.7nC source: http://answers.yahoo.com/question/index?qid=20090702152524AA3IAAr

OpenStudy (anonymous):

okay i want to show how i did it and want to know why im not getting the correct answer-

OpenStudy (anonymous):

\[\rho = q/V\]

OpenStudy (anonymous):

but \[\rho= b/r = 3/2/100 = 300/2\] 300/2 = q/V; V = 4/3 pie [6^3-4^3]=608 *pie/3; 300/2= q/608*pie/3 therefore q = 95456 C..@stormfire1 and @Rohangrr help

OpenStudy (anonymous):

@stormfire1

OpenStudy (vincent-lyon.fr):

@goutham1995 I'll try to explain why you end up with a contradiction. The main reason is this: Your notations are inadequate. The problem is you are calling with the same symbol \(\rho\) different things. You should distinguish: \( \rho\,(r)=\Large \frac{dQ}{dV}=\frac{b}{r}\) which is a function of r, representing how charge-density varies with distance from origin. from: \( \bar \rho=\Large \frac{Q}{V}\) which is the mean value of \(\rho \;(r)\) on your domain. This \( \bar \rho\) can only be worked out when you have calculated Q in the first place. In the light of what I've explained so far, you should understand why your: \(\rho= b/r = 3/2/100 = 300/2\) is 'nonsensical' ;-) What you wrote next, using potential V, has no relation to the question asked. I have not fully checked Rohangrr's answer, but this seems the right method. personally, I do not go into so complicated considerations calculating dV. I write dV = A(r) dr straight away, with A(r) = 4πr², area of the sphere on which the infinitesimal shell is built.

OpenStudy (anonymous):

hey that V which i used was volume not potential

OpenStudy (anonymous):

so \[\rho\] is not equal to 300/2?

OpenStudy (vincent-lyon.fr):

Sorry about my mistake with V for volume! I was mislead by π as in \(\LARGE \frac {1}{4\pi \epsilon _0}\) Which \(\rho\) are you talking about? Where does 300 come from? Where does 2 come from? \(\rho(r)\) varies continuously from \(\rho (R_1)=\Large \frac {b}{R_1}\) to \(\rho (R_2)=\Large \frac {b}{R_2}\)